# Structured Spectral Reasoning for Frequency-Adaptive Multimodal Recommendation

### v2

@@ 3.2 Spectral Band Masking with Consistency Training
Multimodal user–item graphs exhibit frequency components with different semantic roles and unequal reliability ong2024spectrum ; du2023frequency . Low-frequency bands tend to reflect smooth collaborative structure such as long-term preferences or popularity, whereas higher bands capture localized variation that is often modality sensitive. In practice, the latter can correlate with interactions yet fluctuate across items, users and contexts due to sampling noise, content artifacts and instance-specific patterns. Standard training objectives amplify correlations that are easiest to exploit, which biases the predictor toward narrow spectral slices and encourages shortcut behavior. The result is increased sensitivity to small spectral perturbations and weaker generalization under sparsity or cold start. A robustness mechanism should therefore regulate sensitivity at the band level during training while leaving the inference pathway unchanged.
To mitigate this issue, we introduce Spectral Band Masking (SBM), a training-time band-level robustness mechanism. During training, SBM randomly masks a subset of frequency bands and imposes a prediction-consistency objective between the full-spectrum and masked inputs. This structured perturbation reduces over-reliance on any single band, encourages the model to distribute evidence across more stable components and improves robustness under distribution shift. SBM is lightweight, architecture-agnostic and integrates seamlessly with the spectral pipeline.
@@
Let $\{\mathbf{X}^{(m)}\}_{m=1}^{M}$ denote the band-wise features as defined in Section 3.1, and let $\mathbf{x}^{(m)}$ be the per-node slice. During training, we sample a binary mask $\boldsymbol{\gamma}\in\{0,1\}^{M}$ with masking rate $\pi$,
$\gamma_{m}\sim\mathrm{Bernoulli}(1-\pi)\quad(m=1,\dots,M),$ (3)
and construct a masked spectral view
$\tilde{\mathbf{x}}=\sum_{m=1}^{M}\gamma_{m}\,\mathbf{x}^{(m)}.$ (4)
Zeroing entire bands injects a structured perturbation at the spectral level while keeping the graph topology and within-band structure unchanged. We use uniform masking over bands in the main experiments. Energy-aware schedules and curriculum masks can be used as alternatives.
@@
To stabilize training under these spectral perturbations, we pair the masked view with the full-spectrum input $\mathbf{x}=\sum_{m}\mathbf{x}^{(m)}$ and encourage prediction agreement:
$\mathcal{L}_{\text{SBM}}\;=\;\mathbb{E}_{\boldsymbol{\gamma}}\big\|\,f(\mathbf{x})-f(\tilde{\mathbf{x}})\,\big\|_{2}^{2}.$ (5)
This objective acts as band-aware perturbation with a prediction-consistency regularizer. It discourages reliance on any single narrow band, reduces co-adaptation between bands and modalities, and distributes evidence across more stable components. $\mathcal{L}_{\text{SBM}}$ is optimized jointly with the main task loss. The objective adds negligible overhead and makes no changes at inference time.
@@ 3.3 Graph HyperSpectral Neural Operator (G-HSNO)
$\mathbf{Z}=\mathrm{HSNO}(\mathbf{X}),\quad\mathbf{Z}_{v}^{(m)}\;=\;\sum_{n=1}^{M}\mathbf{K}_{mn}\,\mathbf{x}_{v}^{(n)},\quad\text{with }\mathbf{K}_{mn}\in\mathbb{R}^{d\times d}.$ (6)
Collecting $\mathbf{Z}_{v}^{(m)}$ over all nodes yields $\mathbf{Z}\in\mathbb{R}^{|\mathcal{N}|\times M\times d}$. To reduce parameter complexity and avoid overfitting, we adopt a low-rank factorization of the kernel using Canonical Polyadic (CP) decomposition, $\mathbf{K}_{mn}\;=\;\sum_{i=1}^{r}\mathbf{W}_{q}(m,i)\,\mathbf{W}_{k}(n,i)\,\mathbf{V}^{(i)},\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{M\times r},\;\mathbf{V}^{(i)}\in\mathbb{R}^{d\times d},$ where $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ are shared query/key matrices across bands and $r$ is a predefined rank. This structure allows G-HSNO to model second-order interactions between frequency bands while controlling capacity and computation.
@@
To incorporate topological context, we apply a linear projection followed by graph-aware scaling and gating:
$\mathbf{H}_{v}^{(m)}=\sigma\!\big(\mathbf{W}\,\mathbf{Z}_{v}^{(m)}\big)\odot\mathbf{g}_{v}^{(m)},$ (7)
where $\mathbf{W}\in\mathbb{R}^{d^{\prime}\times d}$ is a shared projection, $\sigma(\cdot)$ is a nonlinearity and $\mathbf{g}_{v}^{(m)}\in\mathbb{R}^{d^{\prime}}$ is a learnable frequency-wise gate informed by graph structure (see Section 3.1). We set $d^{\prime}=d$ by default.
The final node representation sums over frequencies, $\mathbf{z}_{v}=\sum_{m=1}^{M}\mathbf{H}_{v}^{(m)}.$ Compared with standard GNN layers or modality-level attention, this step supports structured, spectrum-aware fusion with modest inductive bias. The low-rank kernel controls capacity, and the graph-aware gate enables topologically adaptive adjustments.
@@ 3.4 Spectral Contrastive Regularization
To strengthen semantic alignment and structural integrity in the spectral space, we introduce a lightweight auxiliary objective called Spectral Contrastive Regularization (SCR). Although G-HSNO models frequency-aware graph signals in a structured manner, representations from different modalities can still be inconsistent or collapse within the same band because of noise or weak alignment. Intuitively, components that lie in the same band (for example, image and text embeddings at band $m$) should encode comparable global or local semantics. In practice, distributional heterogeneity can cause divergence or entanglement with unrelated spectral patterns. SCR addresses this by applying a contrastive loss that encourages intra-band cross-modal consistency while preserving separation across bands. Formally, let $\mathbf{x}_{v}^{(m,\mathrm{img})}$ and $\mathbf{x}_{v}^{(m,\mathrm{txt})}$ denote the $m$-th band representations of item $v$ from image and text. These form a positive pair. Negatives are sampled from either different items $v^{\prime}\neq v$ at the same band or from any item at different bands $m^{\prime}\neq m$. The loss is
@@ 3.5 Training and Optimization
The model is trained to predict user–item interaction probabilities from spectral-aware representations. For each pair $(u,v)$, the interaction score is $\hat{y}_{uv}=\sigma(\mathbf{z}_{u}^{\top}\mathbf{z}_{v}),$ where $\mathbf{z}_{u}$ and $\mathbf{z}_{v}$ are the final user and item embeddings, and $\sigma(\cdot)$ is the sigmoid function. The objective combines binary cross-entropy with two auxiliary losses: a band-masking consistency loss $\mathcal{L}_{\text{SBM}}$ that encourages prediction stability under spectral perturbations, and a spectral contrastive loss $\mathcal{L}_{\text{SCR}}$ that aligns modality-specific features within each band. The final loss is
$\mathcal{L}=\mathcal{L}_{\text{BCE}}+\lambda\,\mathcal{L}_{\text{SBM}}+\eta\,\mathcal{L}_{\text{SCR}},$ (9)
where $\lambda,\eta\geq 0$ control the strength of the auxiliary terms. We optimize the model end to end with mini-batch stochastic gradient descent. No masking or rescaling is applied at inference time.
@@ 4.4 Ablation Study and Sensitivity Analysis (RQ3)
We conduct ablations to quantify the contribution of each component: semantic-aware fusion (SAF), the hyperspectral operator (G-HSNO), spectral regularizers (SBM and SCR), and multimodal inputs. As shown in Fig. 2, removing any component leads to clear performance degradation. Notably, removing SAM causes the most significant drop, confirming the importance of semantically guided decomposition. Excluding G-HSNO also harms accuracy, indicating that cross-band and cross-modality reasoning is vital for capturing diverse user–item patterns. SBM and SCR yield smaller but consistent gains, improving robustness and alignment. Eliminating multimodal inputs and using ID features only leads to the weakest results, highlighting the utility of complementary content signals. These findings show that spectral semantics, adaptive fusion and modality reasoning act in a complementary manner within our architecture. We also perform sensitivity analysis over $\lambda$, $\eta$ and $M$, and observe stable performance across a broad range of settings (results are reported in Appendix C.1).
@@ B.4 Implementation Details
We implement our method using PyTorch within the MMRec benchmarking framework [91]. We use 64-dimensional embedding vectors for all item and user representations and apply Xavier initialization [92]. The model is optimized using the Adam optimizer [93] with a batch size of 2048. Learning rates are tuned from $\{0.0001,0.0005,0.001,0.005\}$ using the validation set. Early stopping is triggered if Recall@20 does not improve within 20 validation steps. We conduct extensive ablations on the number of frequency bands $M\in\{2,3,4,5\}$ and evaluate regularization weights for the spectral consistency loss and contrastive term from $\{0.0001,0.001,0.01,0.1,1.0\}$. All hyperparameter choices are selected based on validation performance. Our code and preprocessed datasets will be released to ensure reproducibility.
@@ C.1 Sensitive Analysis
To evaluate the robustness and generalizability of our approach, we conduct parameter sensitivity experiments on three key hyperparameters: the spectral masking regularization weight $\lambda$, the spectral contrastive regularization weight $\eta$, and the number of frequency bands $M$ used in decomposition. As shown in Fig. 2, our model exhibits stable performance across a wide range of values, underscoring its robustness to hyperparameter choices. For $\lambda$, we observe a consistent performance peak around $10^{-2}$, with degradation when $\lambda$ becomes excessively large ($\geq 0.1$), suggesting that overly aggressive suppression of spectral noise may harm useful information flow. A similar pattern emerges for $\eta$, where moderate contrastive regularization ($\eta\approx 10^{-3}\sim 10^{-2}$) improves spectral alignment, but large values quickly deteriorate performance due to over-constraining the representation space. Finally, varying the number of frequency bands $M$ reveals that a three- or four-band decomposition provides the best trade-off: too few bands limit spectral expressiveness, while too many introduce noise and overfitting. These findings validate our design principle of moderate, structure-aware spectral decomposition with targeted regularization, and demonstrate that our model achieves strong results without requiring extensive hyperparameter tuning.

### v3

@@ 3.2 Spectral Band Masking with Consistency Training
Multimodal user–item graphs exhibit frequency components with different semantic roles and unequal reliability Ong and Khong (2024); Du et al. (2023b). Low-frequency bands tend to reflect smooth collaborative structure such as long-term preferences or popularity, whereas higher bands capture localized variation that is often modality sensitive. In practice, the latter can correlate with interactions yet fluctuate across items, users and contexts due to sampling noise, content artifacts and instance-specific patterns. Standard training objectives amplify correlations that are easiest to exploit, which biases the predictor toward narrow spectral slices and encourages shortcut behavior. The result is increased sensitivity to small spectral perturbations and weaker generalization under sparsity or cold start.
To mitigate this issue, we introduce Spectral Band Masking (SBM), a training-time band-level robustness mechanism. During training, SBM randomly masks a subset of frequency bands and imposes a representation-consistency objective between the full-spectrum and masked inputs. This structured perturbation reduces over-reliance on any single band, encourages the model to distribute evidence across more stable components and improves robustness under distribution shift. SBM is lightweight, architecture-agnostic and integrates seamlessly with the spectral pipeline.
@@
Let $\{\mathbf{X}^{(m)}\}_{m=1}^{M}$ denote the band-wise features and let $\mathbf{x}^{(m)}$ be the per-instance slice. During training, we sample a binary mask independently for each instance,
$\gamma^{(i)}_{m}\sim\mathrm{Bernoulli}(1-\pi),\quad m=1,\dots,M,$ (3)
and construct a masked spectral view by dropping entire frequency bands:
$\tilde{\mathbf{x}}^{(i)}=\sum_{m=1}^{M}\gamma^{(i)}_{m}\,\mathbf{x}^{(m)}.$ (4)
In practice, we apply the mask on the band tensor and optionally rescale by the keep probability to keep the expected magnitude stable. To avoid degenerate all-zero masks, we ensure at least one band is kept for each instance. Unless otherwise specified, we use uniform masking over bands.
@@
To stabilize learning under spectral perturbations, we encourage the representation computed from the masked view to match that of the full-spectrum view:
$\mathcal{L}_{\text{SBM}}\;=\;\mathbb{E}_{\boldsymbol{\gamma}}\big\|\,h(\mathbf{x})-h(\tilde{\mathbf{x}})\,\big\|_{2}^{2}.$ (5)
where $\mathbf{h}(\cdot)$ denotes the spectral encoder that outputs representations. This objective acts as band-aware perturbation with a representation-consistency regularizer. It discourages reliance on any single narrow band, reduces co-adaptation between bands and modalities, and distributes evidence across more stable components. $\mathcal{L}_{\text{SBM}}$ is optimized jointly with the main task loss. The objective adds negligible overhead and makes no changes at inference time.
@@ 3.3 Graph HyperSpectral Neural Operator (G-HSNO)
$\mathbf{Z}=\mathrm{HSNO}(\mathbf{X}),\quad\mathbf{Z}_{v}^{(m)}\;=\;\sum_{n=1}^{M}\mathbf{K}_{mn}\,\mathbf{x}_{v}^{(n)}.$ (6)
Collecting $\mathbf{Z}_{v}^{(m)}$ over all nodes yields $\mathbf{Z}\in\mathbb{R}^{|\mathcal{N}|\times M\times d}$. To reduce parameter complexity and avoid overfitting, we parameterize the kernel in a lightweight form. In particular, we adopt a diagonal approximation $\mathbf{K}_{mn}=\mathrm{Diag}(\mathbf{k}_{mn})$, which has been widely used as an efficient spectral filtering variant. Concretely, we factorize $\mathbf{k}_{mn}$ into an inter-band mixing weight and a rank-wise diagonal filter: we compute a band-mixing score $s_{mn}=\langle\mathbf{W}_{q}(m,:),\mathbf{W}_{k}(n,:)\rangle/(\sqrt{r}\tau)$ and normalize it with a softmax function over output bands to $a_{mn}$, and we construct the diagonal filter as $\mathbf{k}_{mn}=a_{mn}\sum_{i=1}^{r}\rho_{i}\ \mathbf{v}^{(i)}$, where $\{\mathbf{v}^{(i)}\}_{i=1}^{r}\subset\mathbb{R}^{d}$ are rank-specific diagonal bases, and $\rho$ is either a global softmax weight or softly adapted by an auxiliary embedding to improve flexibility without introducing a full kernel.
@@
To incorporate instance-specific information into spectral fusion, we learn a lightweight band reweighting function that conditions on the current node representation and its spectral components. Concretely, for node $v$, we compute a band-wise weight vector $\mathbf{g_{v}^{(m)}}$. We then apply the band-wise reweighting to the mixed spectral tensor and aggregate over bands,
$\mathbf{H}_{v}^{(m)}=\mathbf{g}_{v}^{(m)}\,\mathbf{Z}_{v}^{(m)},\qquad\mathbf{z}_{v}=\sum_{m=1}^{M}\mathbf{H}_{v}^{(m)}.$ (7)
To avoid undesired scale drift introduced by reweighting, we further apply a per-instance normalization that aligns the magnitude of the reweighted spectrum with the original one. Compared with fixed band aggregation, this step enables context-adaptive spectral fusion while keeping the overall representation scale stable.
@@ 3.4 Spectral Contrastive Regularization
To strengthen semantic alignment and structural integrity in the spectral space, we introduce a lightweight auxiliary objective called Spectral Contrastive Regularization (SCR). Although G-HSNO models frequency-aware graph signals in a structured manner, representations from different modalities can still be inconsistent or collapse within the same band because of noise or weak alignment. Intuitively, components that lie in the same band (for example, image and text embeddings at band $m$) should encode comparable global or local semantics. In practice, distributional heterogeneity can cause divergence or entanglement with unrelated spectral patterns. SCR addresses this by applying a contrastive loss that encourages intra-band cross-modal consistency while preserving separation across bands. Optionally, the same intra-band contrastive regularization can also be applied to user-side spectral components to further stabilize the spectral space. Formally, let $\mathbf{x}_{v}^{(m,\mathrm{img})}$ and $\mathbf{x}_{v}^{(m,\mathrm{txt})}$ denote the $m$-th band representations of item $v$ from image and text. These form a positive pair. Negatives are sampled from either different items $v^{\prime}\neq v$ at the same band or from any item at different bands $m^{\prime}\neq m$. The loss is
@@ 3.5 Training and Optimization
The model is trained to predict user–item interaction probabilities from spectral-aware representations. For each pair $(u,v)$, the interaction score is $\hat{y}_{uv}=\sigma(\mathbf{z}_{u}^{\top}\mathbf{z}_{v}),$ where $\mathbf{z}_{u}$ and $\mathbf{z}_{v}$ are the final user and item embeddings, and $\sigma(\cdot)$ is the sigmoid function. The objective combines binary cross-entropy with above auxiliary losses: a band-masking consistency loss $\mathcal{L}_{\text{SBM}}$ and a spectral contrastive loss $\mathcal{L}_{\text{SCR}}$ that aligns modality-specific features within each band. Further, we introduce a simple squared regularization on the HSNO kernel and filter to improve training stability. The final loss is
$\mathcal{L}=\mathcal{L}_{\text{BCE}}+\lambda\,\mathcal{L}_{\text{SBM}}+\eta\,\mathcal{L}_{\text{SCR}}+\gamma\,\mathcal{L}_{\text{Reg}},$ (9)
where $\lambda,\eta,\gamma\geq 0$ control the strength of the auxiliary terms. We optimize the model end to end with mini-batch stochastic gradient descent. No masking or rescaling is applied at inference time.
@@ 4.4 Ablation Study and Sensitivity Analysis (RQ3)
We conduct ablations to quantify the contribution of each component: semantic-aware fusion (SAF), the hyperspectral operator (G-HSNO), spectral regularizers (SBM and SCR), and multimodal inputs. As shown in Fig. 2, removing any component leads to clear performance degradation. Notably, removing SAM causes the most significant drop, confirming the importance of semantically guided decomposition. Excluding G-HSNO also harms accuracy, indicating that cross-band and cross-modality reasoning is vital for capturing diverse user–item patterns. SBM and SCR yield smaller but consistent gains, improving robustness and alignment. Eliminating multimodal inputs and using ID features only leads to the weakest results, highlighting the utility of complementary content signals. These findings show that spectral semantics, adaptive fusion and modality reasoning act in a complementary manner within our architecture. We also perform a comprehensive sensitivity analysis over $\lambda$, $\eta$, and $M$, and observe consistently stable performance across a broad and practically relevant range of hyperparameter settings (the detailed results are reported in Appendix C.1).
@@ B.4 Implementation Details
We implement our method using PyTorch within the MMRec benchmarking framework [101]. We use 64-dimensional embedding vectors for all item and user representations and apply Xavier initialization [12]. The model is optimized using the Adam optimizer [26] with a batch size of 2048. Learning rates are tuned from $\{0.0001,0.0005,0.001,0.005\}$ using the validation set. Early stopping is triggered if Recall@20 does not improve within 20 validation steps. We conduct extensive ablations on the number of frequency bands $M\in\{2,3,4,5\}$ and evaluate regularization weights for the spectral consistency loss, contrastive term and regularization term from $\{0.0001,0.001,0.01,0.1,1.0\}$. All hyperparameter choices are selected based on validation performance. Our code and preprocessed datasets will be released to ensure reproducibility.
@@ C.1 Sensitive Analysis
To evaluate the robustness and generalizability of our approach, we conduct parameter sensitivity experiments on three key hyperparameters: the spectral masking regularization weight $\lambda$, the spectral contrastive regularization weight $\eta$, and the number of frequency bands $M$ used in decomposition. As shown in Fig. 2, our model exhibits stable performance across a wide range of values, underscoring its robustness to hyperparameter choices. For $\lambda$, we observe a consistent performance peak around $10^{-2}$, with degradation when $\lambda$ becomes excessively large ($\geq 0.1$), suggesting that overly aggressive suppression of spectral noise may harm useful information flow. A similar pattern emerges for $\eta$, where moderate contrastive regularization improves spectral alignment, but large values quickly deteriorate performance due to over-constraining the representation space. Finally, varying the number of frequency bands $M$ reveals that a three- or four-band decomposition provides the best trade-off: too few bands limit spectral expressiveness, while too many introduce noise and overfitting. These findings validate our design principle of moderate, structure-aware spectral decomposition with targeted regularization, and demonstrate that our model achieves strong results without requiring extensive hyperparameter tuning.